OpenStudy (anonymous):

square root of 3 (5+square root of 3) help me please!

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[sqrt{3} \left( 5+sqrt{3} \right)\]

OpenStudy (dumbcow):

distribute =5sqrt(3) + 3

OpenStudy (anonymous):

i dont get it.

OpenStudy (anonymous):

i wud distribute \[sqrt{3}\] times 5 then square root of 3 times square root of 3 ?

OpenStudy (dumbcow):

yep

OpenStudy (dumbcow):

\[\sqrt{x}*\sqrt{x} = x\]

OpenStudy (anonymous):

i got you now

OpenStudy (dumbcow):

:)